OpenStudy (anonymous):

When f is defined by f(x) = sqrt(x) , find a so f'(a) is five times the value of f'(3).

OpenStudy (anonymous):

hi,

OpenStudy (dumbcow):

\[f'(a) = 5 f'(3)\] First you need to know the function f'(x) \[f'(x) = \frac{d}{dx} \sqrt{x} = \frac{d}{dx} x^{1/2} = \frac{1}{2}x^{-1/2} = \frac{1}{2 \sqrt{x}}\] \[\frac{1}{2\sqrt{a}}=\frac{5}{2\sqrt{3}}\]

OpenStudy (dumbcow):

cancel the 2's , then take reciprocal of both sides \[\sqrt{a} = \frac{\sqrt{3}}{5}\] \[a = \frac{3}{25}\]